OpenStudy (anonymous):

3n+4= 7 Solve the equation , please show your work .

OpenStudy (austinl):

You're solving for n I assume? Subtract 4 from both sides first. Then, divide both sides by 3.

OpenStudy (karategirl2002):

3n means times so 3n+4=7 then n= 1

OpenStudy (anonymous):

3n+4=7 3n=3 n=1

OpenStudy (austinl):

I love how you all just reiterated what I just said in the post above. However there is one difference. I wanted her to do it to learn how to do it herself. Rather than you "show your work" and then just give her the answer. Which is just as bad as giving it outright.

OpenStudy (anonymous):

i have another equation and ts : 11=2-3y

OpenStudy (anonymous):